If x electrons are needed to displace 108g siover from a solution which contains Ag+ ions,how many electrons are needed to displace 9g of aluminium from a solution which contains Al3+ ions?


1
Expert's answer
2021-01-25T04:32:54-0500

According to Faraday's Second Law of electrolysis, when the same quantity of electricity is passed through several electrolytes, the mass of the substances deposited are proportional to their respective chemical equivalent or equivalent weights.

We know silver will undergo the reaction

Ag+ + e- "\to" Ag

Hence, its equivalent weight is "=\dfrac{RAM}{Valence} = \dfrac{108}{1} = 108"

Aluminium on the other hand will deposit as follow;

Al3+ + 3e"\to" Al

and its equivalent weight "= \dfrac{27}{3} = 9"


From Faraday's second law;


"\dfrac{Eq.wt.of Ag}{Eq.wt.of Cu} = \dfrac{Mass. of Ag . deposited}{Mass .of.Cu.depoited}"


"\dfrac{108}{9} = \dfrac{108g}{9g}"


The equation cancels out implying that the number of electrons used in both Ag and Cu is equal = X